# Solutions for $k$-generalized Fibonacci numbers using Fuss-Catalan numbers

### v3

@@ 5 Alternative initial conditions: Spickerman and Joyner
Note that although Spickerman and Joyner [4] gave a correct formal expression for the coefficient $d_{\ell}$, they did not offer an explicit expression for the roots of the auxiliary polynomial. Instead they solved the auxiliary equation numerically for $k=2,\dots,10$ and tabulated the (approximate) numerical values of the roots ([4] Sec. 3). As stated above, our expressions using Fuss-Catalan numbers match their numbers.
6 Multinomial sum
@@ 6.1 General remarks
We digress to consider alternative expressions using multinomial sums. Unlike the previous sections, this is an independent statement of known results. It is included for completeness of the exposition.
@@ 6.2 $k$-generalized Fibonacci
$\begin{split}F_{k}(n+k-1)&=\sum_{i_{1}+2i_{2}+\dots+ki_{k}=n}\frac{(i_{1}+i_{2}+\dots+i_{k})!}{i_{1}!i_{2}!\dots i_{k}!}\\ &=\sum_{i_{1}+2i_{2}+\dots+ki_{k}=n}\binom{i_{1}+i_{2}+\dots+i_{k}}{i_{1},i_{2},\dots,i_{k}}\,.\end{split}$ (6.1)
@@
6.3 More general recurrence
We can generalize the recurrence and multinomial sum as follows. Suppose the recurrence has coefficients $\{\beta_{1},\dots,\beta_{k}\}$, so the recurrence is, omitting too many labels which clutter the notation,
$F(n)=\beta_{1}F(n-1)+\cdots+\beta_{k}F(n-k)\,.$ (6.3)
The generating function is
$G(z;\beta_{1},\dots,\beta_{k})=\frac{z^{k-1}}{1-\beta_{1}z-\dots-\beta_{k}z^{k}}\,.$ (6.4)
Then the multinomial sum is, for $n\geq 0$,
$F(n+k-1)=\sum_{i_{1}+2i_{2}+\dots+ki_{k}=n}\binom{i_{1}+i_{2}+\dots+i_{k}}{i_{1},i_{2},\dots,i_{k}}\,\beta_{1}^{i_{1}}\dots\beta_{k}^{i_{k}}\,.$ (6.5)
Suitable initial values must be specified for $F(0)$ through $F(k-2)$. Note also that the sum in eq. (6.5) is normalized to yield $F(k-1)=1$.
@@ 6.4 Narayana sequence
$N(n)=N(n-1)+N(n-3)\,.$ (6.6)
Hence we set $k=3$ and $(\beta_{1},\beta_{2},\beta_{3})=(1,0,1)$ in Sec. 6.3. Because $\beta_{2}=0$, we exclude all terms in $i_{2}$ in eq. (6.5) with $i_{2}>0$. This yields, using $i_{1}=n-3i_{3}$,
$F(n+2)=\sum_{i_{1}+3i_{3}=n}\binom{i_{1}+i_{3}}{i_{1},i_{3}}=\sum_{i_{3}=0}^{\lfloor(n/3)\rfloor}\binom{n-2i_{3}}{i_{3}}\,.$ (6.7)
Set the initial values $F(0)=F(1)=1$, then $F(n)=N(n)$. The first few terms are $1,1,1,2,3,4,6,9,13,19,\dots$.
@@ 6.5 Padovan sequence
$P(n)=P(n-2)+P(n-3)\,.$ (6.8)
Hence we set $k=3$ and $(\beta_{1},\beta_{2},\beta_{3})=(0,1,1)$ in Sec. 6.3. Because $\beta_{1}=0$, we exclude all terms in $i_{1}$ in eq. (6.5) with $i_{1}>0$. This yields
$F(n+2)=\sum_{2i_{2}+3i_{3}=n}\binom{i_{2}+i_{3}}{i_{2},i_{3}}\,.$ (6.9)
With the initial values $F(0)=F(1)=0$, this yields the sequence $0,0,1,1,1,2,2,3,4,5,7,9,12,16,\dots$. Hence $P(n)=F(n+2)$.

### v4

@@ 6 Alternative initial conditions: Spickerman and Joyner
Note that although Spickerman and Joyner [11] gave a correct formal expression for the coefficient $d_{j}$, they did not offer an explicit expression for the roots of the auxiliary polynomial. Instead they solved the auxiliary equation numerically for $k=2,\dots,10$ and tabulated the (approximate) numerical values of the roots ([11] Sec. 3). As stated above, our expressions using Fuss-Catalan numbers match their numbers.
7 Basis sequences
7.1 $k$-generalized Fibonacci
The $k$-generalized Fibonacci numbers $F_{k}(n)$ are the solution of the recurrence eq. (1.1) for only one set of initial conditions, viz. $F_{k}(0)=\dots=F_{k}(k-2)=0$ and $F_{k}(k-1)=1$. We would like a set of $k$ ‘basis sequences’ or ‘fundamental solutions’ $\{B_{k,i}(n),\,i=0,\dots,k-1\}$ of the recurrence eq. (1.1), where the sequence $B_{k,i}(n)$ satisfies the initial conditions $B_{k,i}(n)=\delta_{in}$ for $n=0,\dots,k-1$. The $k$-generalized Fibonacci numbers $F_{k}(n)$ are the last member of the basis, viz. $F_{k}(n)=B_{k,k-1}(n)$.
Wolfram [13] solved the problem in a more general setting. Wolfram treated the homogeneous linear recurrence ([13], eq. (2.1))
$f(n)=\sum_{j=1}^{k}a_{k-j}f(n-j)\,.$ (7.1)
Here $\{a_{0},\dots,a_{k-1}\}$ are a set of coefficients which do not depend on $f$ but can depend on $n$. Let us define $\mathscr{W}_{i}(n)$ to be the solution of eq. (7.1) with the initial values $\mathscr{W}_{i}(n)=\delta_{in}$ for $n=0,\dots,k-1$. (For brevity, we omit explicit mention of $k$ and $a_{0},\dots,a_{k-1}$.) The set of sequences $\{\mathscr{W}_{i}(n),\,i=0,\dots,k-1\}$ are a set of fundamental solutions of the recurrence eq. (7.1). Wolfram proved that ([13] Lemma 2.2)
$\mathscr{W}_{i}(n)=\sum_{j=0}^{i}a_{i-j}\mathscr{W}_{k-1}(n-j-1)\qquad\qquad(i=0,\dots,k-2)\,.$ (7.2)
Hence it is only necessary to solve the recurrence eq. (7.3) for the last basis function $\mathscr{W}_{k-1}(n)$. All the other basis functions can be obtained from it. This is a key insight.
We treat only constant coefficients in this paper. Let $\{\beta_{1},\dots,\beta_{k}\}$ be a set of constant coefficients and reexpress eq. (7.1) as follows. (Note the indexing of the $\beta_{j}$, to maintain consistency with other usage of indexing in this paper.)
$f(n)=\sum_{j=1}^{k}\beta_{j}f(n-j)\,.$ (7.3)
We define $W_{i}(n)$ to be the solution of eq. (7.3) with the initial values $W_{i}(n)=\delta_{in}$ for $n=0,\dots,k-1$. Then, from eq. (7.2),
$W_{i}(n)=\sum_{s=0}^{i}\beta_{k+s-i}W_{k-1}(n-s-1)\qquad\qquad(i=0,\dots,k-2)\,.$ (7.4)
We shall employ $W_{i}(n)$ and eq. (7.4) in the rest of this paper.
Proposition 7.1.
The basis sequences for the $k$-generalized Fibonacci numbers are given by
$B_{k,k-m}(n)=\sum_{j=0}^{k-1}\frac{\zeta_{j}^{m}-2\zeta_{j}^{m-1}+1}{(k+1)\zeta_{j}-2k}\zeta_{j}^{n-k+1}\qquad(m=1,\dots,k)\,.$ (7.5)
Proof.
The $k$-generalized Fibonacci recurrence eq. (1.1) is the special case $\beta_{1}=\dots=\beta_{k}=1$ of eq. (7.3). It follows immediately from eq. (7.4), where our last basis function is $F_{k}(n)$, that
$B_{k,i}(n)=\sum_{s=0}^{i}F_{k}(n-s-1)\,.$ (7.6)
We employ the expression in eq. (4.7) and set i=kmi=k-m. In the line tagged by ()(*) below, we employ the auxiliary equation to deduce $\zeta_{j}^{-k-1}=2\zeta_{j}^{-1}-1$, because $\zeta_{j}^{k+1}-2\zeta_{j}^{k}+1=0$ (see eq. (2.12)).
$\begin{split}B_{k,k-m}(n)&=\sum_{s=0}^{k-m}F_{k}(n-s-1)\\ &=\sum_{s=0}^{k-m}\sum_{j=0}^{k-1}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\frac{\zeta_{j}^{n-k+1}}{\zeta_{j}^{s+1}}\\ &=\sum_{j=0}^{k-1}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\zeta_{j}^{n-k+1}\biggl{(}\frac{1}{\zeta_{j}}\sum_{s=0}^{k-m}\frac{1}{\zeta_{j}^{s}}\biggr{)}\\ &=\sum_{j=0}^{k-1}\frac{\zeta_{j}-1}{(k+1)\zeta_{j}-2k}\zeta_{j}^{n-k+1}\frac{1-\zeta_{j}^{m-k-1}}{\zeta_{j}-1}\\ &=\sum_{j=0}^{k-1}\frac{1}{(k+1)\zeta_{j}-2k}\zeta_{j}^{n-k+1}(1-\zeta_{j}^{m}(2\zeta_{j}^{-1}-1))\qquad\qquad(*)\\ &=\sum_{j=0}^{k-1}\frac{1}{(k+1)\zeta_{j}-2k}\zeta_{j}^{n-k+1}(\zeta_{j}^{m}-2\zeta_{j}^{m-1}+1)\,.\end{split}$ (7.7)
Rearranging terms in the numerator yields eq. (7.5). ∎
Remark 7.2.
Observe that eq. (7.5) is a Binet-style sum for each value of $m$, and is valid for negative $n$. The dependence on $m$ has a simple pattern. The case $m=1$ equals eq. (4.7). To the author’s knowledge, eq. (7.5) has not been published in the literature.
Next, given a tuple of initial values $\vec{\gamma}=(\gamma_{0},\dots,\gamma_{k-1})$ for $n=0,\dots,k-1$, the solution $B(n,k,\vec{\gamma})$ of the recurrence eq. (1.1) is given by the linear combination
$B(n,k,\vec{\gamma})=\sum_{i=0}^{k-1}\gamma_{i}B_{k,i}(n)\,.$ (7.8)
We can leave it at this, because eq. (7.5) furnishes expressions for all the basis functions $B_{k,i}(n)$.
However, Wolfram’s formalism [13] and eq. (7.4) furnishes tools to offer a solution for the more general recurrence in eq. (7.3). Using eq. (7.4), we can also express everything in terms of only the last basis function, which is $W_{k-1}(n)$. Given a tuple of initial values $\vec{\gamma}=(\gamma_{0},\dots,\gamma_{k-1})$ for $n=0,\dots,k-1$, denote the solution of the recurrence eq. (7.3) by $W(n,k,\vec{\gamma})$. Then
$\begin{split}W(n,k,\vec{\gamma})&=\sum_{i=0}^{k-1}\gamma_{i}W_{i}(n)\\ &=\gamma_{0}\beta_{k}W_{k-1}(n-1)\\ &\quad+\gamma_{1}\bigl{(}\beta_{k-1}W_{k-1}(n-1)+\beta_{k}W_{k-1}(n-2)\bigr{)}\\ &\quad+\dots\\ &\quad+\gamma_{k-2}\bigl{(}\beta_{2}W_{k-1}(n-1)+\beta_{3}W_{k-1}(n-2)+\dots+\beta_{k}W_{k-1}(n-k+1)\bigr{)}\\ &\quad+\gamma_{k-1}W_{k-1}(n)\\ &=\bigl{(}\beta_{k}\gamma_{0}+\beta_{k-1}\gamma_{1}+\dots+\beta_{2}\gamma_{k-2}\bigr{)}W_{k-1}(n-1)\\ &\quad+\bigl{(}\beta_{k}\gamma_{1}+\beta_{k-1}\gamma_{2}+\dots+\beta_{3}\gamma_{k-2}\bigr{)}W_{k-1}(n-2)\\ &\quad+\dots\\ &\quad+\beta_{k}\gamma_{k-2}W_{k-1}(n-k+1)\\ &\quad+\gamma_{k-1}W_{k-1}(n)\\ &=\gamma_{k-1}W_{k-1}(n)+\sum_{i=1}^{k-1}W_{k-1}(n-i)\biggl{(}\sum_{j=2}^{k-i+1}\beta_{j+i-1}\gamma_{k-j}\biggr{)}\,.\end{split}$ (7.9)
The nested sums over $j$ do not depend on $n$, hence they can be precomputed, say $\delta_{i}\equiv\sum_{j=2}^{k-i+1}\beta_{j+i-1}\gamma_{k-j}$ for $i=1,\dots,k-1$ and formally define $\delta_{0}=\gamma_{k-1}$. Then
$W(n,k,\vec{\gamma})=\sum_{i=0}^{k-1}\delta_{i}W_{k-1}(n-i)\,.$ (7.10)
8 Companion matrix
This section was prompted by an insightful comment by a referee. It contains alternative formalism for the solution of linear recurrences with constant coefficients. Our exposition follows Chen and Louck [2]. Recall the recurrence in eq. (7.3), reproduced here for ease of reference.
$f(n)=\beta_{1}f(n-1)+\cdots+\beta_{k}f(n-k)\,.$ (8.1)
The characteristic polynomial of the recurrence in eq. (8.1) is $p_{k}(x)=x^{k}-\beta_{1}x^{k-1}-\beta_{2}x^{k-2}-\dots-\beta_{k}$. The companion matrix $C_{k}$ of $p_{k}(x)$ is defined as follows ([2], eq. (2.1))
$C_{k}=\begin{pmatrix}\beta_{1}&\beta_{2}&\beta_{3}&\dots&\beta_{k-1}&\beta_{k}\\ 1&0&0&\dots&0&0\\ 0&1&0&\dots&0&0\\ \vdots&\vdots&\vdots&\dots&\vdots&\vdots\\ 0&0&0&\dots&1&0\end{pmatrix}\,$. (8.2)
The characteristic equation of the matrix $C_{k}$ is $p_{k}(\lambda)=0$, which is also the characteristic equation of the recurrence in eq. (8.1). (For the special case $\beta_{1}=\dots=\beta_{k}=1$, the Fuss-Catalan series for the roots $\zeta_{j}$, $j=0,\dots,k-1$ in eqs. (2.7) and (2.8) are the eigenvalues of the companion matrix $C_{k}$.) Then ([2] Theorem 3.1) the $(i,j)$ entry $c_{i,j}^{(n)}$ in the matrix $C_{k}^{n}$ is given by ([2] eq. (3.1), note that the indexing is $1\leq i,j\leq k$)
$c_{i,j}^{(n)}=\sum_{i_{1}+2i_{2}+\dots+ki_{k}=n-i+j}\frac{i_{j}+\dots+i_{k}}{i_{1}+\dots+i_{k}}\binom{i_{1}+i_{2}+\dots+i_{k}}{i_{1},i_{2},\dots,i_{k}}\,\beta_{1}^{i_{1}}\dots\beta_{k}^{i_{k}}\,.$ (8.3)
Also, $c_{i,j}^{(n)}$ is defined to equal unity if n=ijn=i-j. The companion matrix can be employed to iterate the solution of a linear recurrence with constant coefficients as follows ([2], Sec. 7). Iterating $n\geq 1$ times yields ([2] eq. (7.3))
$\begin{pmatrix}f(n+k-1)\\ f(n+k-2)\\ \vdots\\ f(n)\end{pmatrix}=C_{k}^{n}\,\begin{pmatrix}f(k-1)\\ f(k-2)\\ \vdots\\ f(0)\end{pmatrix}\,$. (8.4)
The value of f(n+ki)f(n+k-i) in left-hand column vector is given by the dot product of the $i^{th}$ row of the matrix $C_{k}^{n}$ with the $k$-tuple of initial values $f(0)$ through f(k1)f(k-1). For fixed $1\leq j\leq k$, if the initial $k$-tuple is $(f(0),\dots,f(k-1))=(\dots,\delta_{ij},\dots)$, where $\dots$” denotes zeroes, the solution for f(n+ki)f(n+k-i) is given by $c_{i,j}^{(n)}$, $1\leq i\leq k$. It is evident that the matrix elements in the $j^{th}$ column, $1\leq j\leq k$, are the elements of the $(k-j)^{th}$ basis sequence of the recurrence, viz. $W_{k-j}(n-k-i)=c_{i,j}^{(n)}$, for $1\leq i\leq k$.
Proposition 8.1.
Wolfram’s Lemma 2.2 [13] can be employed to express $c_{i,j}^{(n)}$ for $j=2,\dots,k$ in terms of the entries $c_{i,1}^{(\dots)}$ in the first column of the companion matrix as follows
$c_{i,j}^{(n)}=\sum_{s=0}^{\min(k-j,n-1)}\beta_{j+s}c_{i,1}^{(n-s-1)}\qquad\qquad(j=2,\dots,k)\,.$ (8.5)
Proof.
Transcribe the notation in eq. (7.4), accounting for differences in the indexing. ∎
Hence for $n>1$ it is only necessary to compute the entries in the first column of the companion matrix from scratch; all the rest are linear combinations of those numbers. This may be computationally helpful. Eq. (8.5) has been numerically tested and validated. Wolfram’s paper [13] was published in 2000, i.e. recently, hence the above identity may not be known in the literature for companion matrices. It is not mentioned by Chen and Louck [2], whose paper is dated 1996.
9 Generating functions and multinomial sums
@@ 9.1 General remarks
We remark on alternative expressions using generating functions and multinomial sums. A referee kindly brought to the author’s attention that the multinomial sum in eq. (9.4) below is a Dickson polynomial of the second kind in several variables. The connection to Dickson polynomials is shown below.
  1. 1.
    In this section, the arguments of all functions are nonnegative integers. One can run all the recurrences backwards, but the multinomial sums displayed below are valid only for nonnegative arguments.
  2. 2.
    To avoid ambiguity, we apply a circumflex, e.g. $\hat{f}$, to denote the solution of a recurrence when the initial $k$-tuple is $(0,\dots,0,1)$.
@@ 9.2 $k$-generalized Fibonacci
[…]
@@
9.3 Linear recurrence with constant coefficients
We can generalize the recurrence and multinomial sum as follows. We employ the recurrence in eq. (7.3). The generating function is
$G(z)=\frac{z^{k-1}}{1-\beta_{1}z-\dots-\beta_{k}z^{k}}\,.$ (9.3)
By construction, the initial values are $f(0)=\dots=f(k-2)=0$ and f(k1)=1f(k-1)=1 and the remaining values of are given by the multinomial sum, for $n\geq 0$,
$\hat{f}(n+k-1)=\sum_{i_{1}+2i_{2}+\dots+ki_{k}=n}\binom{i_{1}+i_{2}+\dots+i_{k}}{i_{1},i_{2},\dots,i_{k}}\,\beta_{1}^{i_{1}}\dots\beta_{k}^{i_{k}}\,.$ (9.4)
Remark 9.1.
Levesque [8] published a similar generating function and multinomial sums. However, the numerator of Levesque’s generating function incorporated arbitrary initial values for $f(0),\dots,f(k-1)$. Our precomputed sums of linear combinations of the initial values in Sec. 7 are similar to Levesque’s expressions for $v_{i}$ in ([8] eq. (2.2)).
A referee kindly brought to the author’s attention that the sum in eq. (9.4) is a Dickson polynomial of the second kind in $k$ variables. (The Dickson polynomials of the first and second kind, in one variable, are related to the Chebyshev polynomials.) Observe that $\hat{f}(n+k-1)$ in eq. (9.4) equals the matrix element $c_{1,1}^{(n)}$ (see eq. (8.3)). The Dickson polynomial of the second kind in $k$ variables also equals $c_{1,1}^{(n)}$. See eq. (5.2) by Chen and Louck [2], with the transcriptions $m\leftarrow k-1$ and $u_{i}\leftarrow\beta_{i}$. From Prop. 8.1, all the other matrix elements $c_{i,j}^{(n)}$ are linear combinations of Dickson polynomials.
@@ 9.4 Narayana sequence
N(n)=N(n1)+N(n3).N(n)=N(n-1)+N(n-3)\,. (9.5)
Then $k=3$ in eq. (8.1) and the coefficients are $(\beta_{1},\beta_{2},\beta_{3})=(1,0,1)$. We employ the basis functions $W_{i}(n)$ from Sec. 7 for this and the other sequences treated below. The basis function with the initial values $(0,0,1)$ is $W^{N}_{2}(n)$. (We attach a superscript ‘N’ for Narayana.) The other basis functions are $W^{N}_{0}(n)=\beta_{3}W^{N}_{2}(n-1)=W^{N}_{2}(n-1)$ and $W^{N}_{1}(n)=\beta_{2}W^{N}_{2}(n-1)+\beta_{3}W^{N}_{2}(n-2)=W^{N}_{2}(n-2)$. Using the given initial values, the solution is
$\begin{split}N(n)&=W^{N}_{0}(n)+W^{N}_{1}(n)+W^{N}_{2}(n)\\ &=W^{N}_{2}(n-1)+\underbrace{W^{N}_{2}(n-2)+W^{N}_{2}(n)}_{=W^{N}_{2}(n+1)}\\ &=W^{N}_{2}(n-1)+W^{N}_{2}(n+1)\\ &=W^{N}_{2}(n+2)\,.\end{split}$ (9.6)
For the multinomial sum, because $\beta_{2}=0$, we exclude all terms in $i_{2}$ in eq. (9.4) with $i_{2}>0$. This yields, using $i_{1}=n-3i_{3}$,
$W^{N}_{2}(n+2)=\sum_{i_{1}+3i_{3}=n}\binom{i_{1}+i_{3}}{i_{1},i_{3}}=\sum_{i_{3}=0}^{\lfloor(n/3)\rfloor}\binom{n-2i_{3}}{i_{3}}\,$. (9.7)
The first few values of $N(n)$ are $1,1,1,2,3,4,6,9,13,19,\dots$.
@@ 9.5 Padovan sequence
P(n)=P(n2)+P(n3).P(n)=P(n-2)+P(n-3)\,. (9.8)
Then $k=3$ in eq. (8.1) and the coefficients are $(\beta_{1},\beta_{2},\beta_{3})=(0,1,1)$. The basis function with the initial values $(0,0,1)$ is $W^{P}_{2}(n)$. (We attach a superscript ‘P’ for Padovan, and it can also mean Perrin in Sec. 9.6 below.) The other basis functions are $W^{P}_{0}(n)=\beta_{3}W^{P}_{2}(n-1)=W^{P}_{2}(n-1)$ and $W^{P}_{1}(n)=\beta_{2}W^{P}_{2}(n-1)+\beta_{3}W^{P}_{2}(n-2)=W^{P}_{2}(n-1)+W^{P}_{2}(n-2)$. Using the given initial values, the solution is
$\begin{split}P(n)&=W^{P}_{0}(n)+W^{P}_{1}(n)+W^{P}_{2}(n)\\ &=\underbrace{W^{P}_{2}(n-1)+W^{P}_{2}(n-2)}_{=W^{P}_{2}(n+1)}+\underbrace{W^{P}_{2}(n-1)+W^{P}_{2}(n)}_{=W^{P}_{2}(n+2)}\\ &=W^{P}_{2}(n+1)+W^{P}_{2}(n+2)\\ &=W^{P}_{2}(n+4)\,.\end{split}$ (9.9)
For the multinomial sum, because $\beta_{1}=0$, we exclude all terms in $i_{1}$ in eq. (9.4) with $i_{1}>0$. This yields
$W^{P}_{2}(n+2)=\sum_{2i_{2}+3i_{3}=n}\binom{i_{2}+i_{3}}{i_{2},i_{3}}\,$. (9.10)
The first few values of $P(n)$ are $1,1,1,2,2,3,4,5,7,9,12,16,\dots$.